368070:

Which function represents the area of the rectangle? 6x2 – 8x + 36 6x2 – 27x + 36 6x2 – 35x + 36 6x2 + 35x + 36

TutoriaIs:

Which Rectangle?

368070:

1 attachment
Fan999:

Is there any more information?

Fan999:

Nvm ^

368070:

yeah I just put in a screenshot

368070:

oopps never mind too

Fan999:

Did you figure it out?

368070:

nope

Fan999:

Okay, do you know how to find the function? Or find A function?

368070:

I used to but I forgot sorry im dumb hahaha

Fan999:

No problem I can help :)

368070:

okay thanks you a lifesaver

snowflake0531:

You want to find \((3x-4)(2x-9)\) and you can use the foil method \((a+b)(c+d) = ac+ad+bc+bd\) As it is you can do \( 3x(2x) + 3x(-9) + (-4)(2x) + (-4)(-9)\) Can you simplify the terms?

Fan999:

https://www.youtube.com/watch?v=JySCLnLV8hY Use this video to help a little better.

368070:

thank you

368070:

yup i'm still lost

368070:

someone need to dumb this down as much as they can

Fan999:

Alright... let me try

Fan999:

So you want to find (3x-4) And (2x-9) To do so you can do (a+b) (c+d)= ac+ad+bc+bd Plug in your numbers from your equation So that (^) becomes: 3x(2x)+3x(−9)+(−4)(2x)+(−4)(−9) Now simplify that. Does that make any more sense? Or are you still confused...

Fan999:

After you simplify you should get one of your answer choices.

snowflake0531:

love how you copy me without changing anything <3

NotLeah:

Multiply each term in the first expression by each term in the second expression and combine like terms. 6x2 – 35x + 36

368070:

thank you girlll

Fan999:

@368070 wrote:
thank you girlll
Sure thing, glad you figured it out!